# To FP8 and Back Again: Quantifying the Effects of Reducing Precision on LLM Training Stability

### v1

@@ 3.2 Masking
Figure 3: Diagram showing the precisions used in a Llama decoder block (best seen in color). The activations in the path of the residual connection are kept in FP32, as are the model weights and embeddings. The LayerNorm and RoPE layers use FP32 internally for their computations. The Flash Attention kernel uses BF16 with no reduction in precision. All other layers use reduced-precision matrix multiplication that emulates low-precision computation with a high-precision accumulator.
Algorithm 1 PyTorch-like pseudocode for the reduced-precision forward pass.
def forward(input, weight):
save_for_backward(input, weight)
masked_input = reduce_precision(input)
masked_weight = reduce_precision(weight)
output = F.linear(masked_input, masked_weight)
masked_output = reduce_precision(output)
return masked_output
Figure 4: Exponent masking implemented by clamping values that cannot be expressed with the allowed number of exponent bits.
Our experiments use a simplified method of reducing floating-point precision to achieve reasonable throughput. To simulate removing exponent bits, we threshold the values to the minimum and maximum absolute values possible with the given number of exponent bits. Figure 4 depicts the exponent masking process. To remove the mantissa bits, a bitmask is applied to remove the unrepresentable bits. The resulting method is an imperfect approximation of reducing the bit count of floating-point numbers. However, it has the advantage of being fast at run time, causing at most a doubling of the time per each training step.
Reduced precision operations are applied only on the matrix multiplication computations of the model, excluding the attention computation, which uses the Flash Attention v2 kernel. Following existing FP8 libraries such as TransformerEngine [32], we separate the effects of reducing the computation’s precision from reducing the data’s precision in storage. As a result, the activations and model weights are kept at their original precision while the inputs and outputs of matrix multiplication are dynamically masked to emulate reduced precision computation with a high-precision accumulator. All states are kept in their original precision, and all operations other than matrix multiplication are performed in their original precision. In Figure 3, we include a diagram indicating the precision of the states and computations in a Llama decoder block.
@@ 4.1 MS-AMP experiments
We first analyze the effect of real-world FP8 training by applying the MS-AMP [10] library (version 0.3.0) to the nanoGPT codebase. We run all experiments for MS-AMP on an H100 node with 8 GPUs to ensure hardware availability of FP8. However, despite only using the O1 optimization level for MS-AMP, we find that the resulting models show non-trivial performance degradations. These results indicate either that MS-AMP cannot converge to the same loss as BF16 training or that it requires more training steps. In either case, the justification for deploying FP8 training as a cost-saving measure becomes weaker.
Figure 5: Comparison between training with BF16 and with MS-AMP level O1. Experiments on MS-AMP were conducted on the nanoGPT codebase with only the model architecture as a variable. In both cases, even level O1 optimization of MS-AMP underperforms BF16 training noticeably. We use a sequence length of 1024 and a global batch size of 480 on a single node with eight H100 GPUs. The nanoGPT codebase was used because we were unable to integrate TinyLlama with MS-AMP.
@@ 4.2 Bit reduction experiments
Figure 6: TinyLlama 120M models trained until loss divergence occurred. E8M3, E8M4, and E8M5 models trained for 16K, 20K, and 100K steps, respectively. The dotted black line in each figure indicates the loss landscape sharpness of the model. A global batch size of 512 was used with a sequence length of 2048.
Figure 7: Llama 7B model training loss curves for different mantissa bit masks. The x-axis represents training steps, while the y-axis represents the training loss.
@@
We found that removing even a single exponent bit prevents training altogether, resulting in the model failing to progress with any learning using E7M7, confirming previous findings [24] that neural network training is more sensitive to exponent bits than mantissa bits. To analyze the cause, we first conduct an ablation on the clamping mechanism by either removing only the inner or outer exponent range, as depicted in Figure 4. We find that models with only their inner exponent ranges clamped train normally while models with only their outer exponent ranges clamped could not engage in training, indicating that the inability to represent large values is the cause of failure for E7M7.
We also found that specific regions of the model are more sensitive to reduced precision than others. For the Llama models, removing the masking from the LM head was sufficient to restore training, although the resulting model was much more unstable. This result supports the conclusion that certain parts of the model are more sensitive to reduced precision than others, as no training occurs when the entire model is trained with E7M7.
4.3 Loss landscape sharpness
To further uncover the relationship between bit width and training robustness, we use Equation 2 to quantify the degree of training instability increase by measuring the loss landscape instability of Llama 7B models in Table 1 and Llama 120M models in Figure 6.
In Figure 6, we show Llama 120M models trained until training loss divergence occurs, as well as plotting the loss landscape sharpness values of the models in E8M3, E8M4, and E8M5. Although the points of divergence are different for each of the models, we can see a general trend of increasing sharpness until the model diverges sharply, after which it cannot revert to its original training trajectory. Training required approximately 2 hours per 10,000 training steps on a node with eight A100 GPUs.
To verify that similar behavior occurs in larger models, we include a comparison of the training losses of Llama v2 models with 7B parameters trained for 5,000 steps in Figure 7. We apply early stopping at 5,000 training steps because training a Llama 7B model for 5K steps requires approximately one week on a node with eight A100 GPUs. From these experiments, we can see that loss divergence is visible in the E8M3 and E8M4 models, while it has yet to emerge in the E8M5 model. However, from Table 1, we can see that the loss landscape sharpness continues to increase for the E8M5 model, even though no signs of instability are yet visible.
The E8M3 and E8M4 models show much higher sharpness values and both accordingly diverge early in training. In contrast, there is only a gradual increase in the loss-landscape sharpness for the E8M7 runs. As seen in Figures 7 and 6, we observe that models gradually increase in sharpness until a threshold is reached, although the threshold value may differ depending on the configurations. These results suggest that models with fewer mantissa bits are entering regions of ever greater instability during training, even when these instabilities are not visible in the loss curve. We believe that, in the future, such analysis of loss landscape sharpness can be used to identify when the model is at risk of loss divergence.
@@ 4.4 Robustness to learning rate changes
We further attempt to identify hidden instability in E8M5, which did not diverge during the initial training stages in Figure 6. Inspired by Wortsman et al. [18], we analyze the robustness of Llama 120M models to changes in the learning rate by comparing training at BF16 with that for E8M5. As seen in Figure 8, the E8M5 training runs have more frequent loss spikes during training, especially when the learning rate is increased to $4e\mathrm{-3}$. Although no cases of loss divergence were found, we believe that the higher frequency of loss spikes indicates greater instability during training, supporting our claim that the loss landscape is sharper for E8M5 even before loss divergence occurs.
@@ 5 Discussion
This work aims to propose quantitative evaluations and analyses of training instabilities that occur when reducing floating-point precision. By clamping unrepresentable exponent values and masking out the mantissa before the computation, we can approximate matrix multiplication with reduced-precision computation and high-precision accumulation. Our experiments have shown that reducing the precision of floating-point computations destabilizes LLM training and that existing mechanisms to stabilize FP8 training do not offer sufficient robustness as to allow their cost-effective use.
Again, we emphasize that our work does not aim to disprove the viability of FP8 training. We merely argue that FP8 training, in its current form, is uneconomical, not impractical. Even if FP8 training is effective for carefully selected hyperparameters under specific conditions, our assertion is that the costs of finding such conditions and the risks involved in training a less stable model outweigh the benefits of using FP8 for accelerated computation. Instead, we propose methods to evaluate the stability and robustness of reduced-precision training, which is key for FP8 or other reduced-precision training schemes to be viable for LLM training in the real world.
Also, we would like to preempt misunderstanding by making clear that we are not questioning the usefulness of FP8 for inference. Several recent works [33, 34, 35] have shown that it is even possible to quantize LLM weights to below 4 bits without sacrificing much accuracy. Xia et al. [36] has also shown that, even for the A100, using FP6 for inference is a viable option. We believe that dedicated FP8 and FP4 processors can make LLM inference simpler to implement and faster to compute.
We note that, from our experiments, several methods naturally suggest themselves as possible stabilization techniques. First, the initial stages of training could be conducted in higher precision, similar to how smaller batch sizes may be used during the initial stages of training [20]. Increasing the precision when the loss landscape becomes too sharp and returning to lower precision when the model is in a smoother region may also provide a tradeoff between training speed and stability.
Second, the layers that are more sensitive to reduced precision may be kept at high precision, while only the less sensitive model layers are computed with reduced precision. For example, during our experiments, we found that removing masking from the LM head of a Llama model was sufficient to enable E7M7 training, although the resulting model was less stable. For GPT models, we found that increasing the precision of the first two decoder blocks to TF32 was sufficient to prevent loss divergence. However, as such compensatory techniques depend on the model architecture, training data, and other aspects of the training environment, we leave their investigation to future work.
@@ 6 Limitations
A limitation of this work is that it focuses on the initial stages of pre-training when many instabilities are known to arise only later in training [37]. For example, Wortsman et al. [18] show that the logits of the outputs diverge from zero only at the later stages of training. To this, we argue that our studies likely underestimate the instabilities that FP8 or other reduced precision training schemes will face, further strengthening our case that reduced-precision training methods are too unstable to be profitably utilized in their current form.
Second, despite finding that the exponent bits are of greater importance to LLM training than mantissa bits, we were unable to experiment by increasing the number of exponent bits. This was because matrix multiplication in FP64 is over an order of magnitude slower than BF16 on A100 and H100 GPUs when using tensor cores. Experiments using representations such as E11M4, created by removing 48 mantissa bits from FP64, may be illuminating, but we found it impractical to train models with a greater number of exponent bits.
Finally, our experiments are limited in that they only the training loss is used as an evaluation metric instead of real-world natural language tasks such as MMLU [38] scores. However, while lower perplexity is no guarantee of superior performance on downstream tasks, we believe that the divergence of the training loss is sufficient as an indicator of training failure.
7 Societal impact
As the costs of training or even fine-tuning LLMs continue to climb with the ever-growing size of the largest models and datasets, we believe that FP8 can be a useful tool in reducing the costs of LLMs to manageable levels. The recent introduction of FP8 in the NVIDIA Ada and Hopper [39] series of GPUs and the even more recent announcement of FP4 support in the Blackwell [40] series of GPUs is placing these low-precision data types within the reach of industry and academia, both of which groups are always in need of ever more compute.
Although we have focused on the difficulties of using FP8, with the electricity consumption of data centers projected to double within the next few years [41], mostly due to increased artificial intelligence workloads, we believe that any improvement in the training efficiency of LLMs would have a noticeable impact on global electricity consumption. As such, it is our hope that future work uses our work as a guide to improving the ways in which low-precision computations can be used.
@@ 8 Conclusion
We demonstrate that the training stability of LLMs decreases incrementally with the reduction of floating-point bit widths used for training models of up to 7B parameters. Using our proposed loss landscape sharpness metric, we measure the gradual increase of instability that leads to loss divergence, shedding light on a phenomenon with potentially large financial and environmental costs.

### v2

@@ 3.2 Masking
Our experiments use a simplified method of reducing floating-point precision to achieve reasonable throughput. To simulate removing exponent bits, we threshold the values to the minimum and maximum absolute values possible with the given number of exponent bits. Figure 10 depicts the exponent masking process. A bitmask is applied to remove the unrepresentable mantissa bits. The resulting method is an imperfect approximation of reducing floating-point bit count. However, it has the advantage of being fast, causing at most a doubling of the time per training step.
We apply reduced precision operations only on the matrix multiplication computations of the model, excluding the attention computation, which uses the Flash Attention v2 kernel. Following existing FP8 libraries such as TransformerEngine (NVIDIA, 2023), we separate the effects of reducing the computation’s precision from reducing the data’s precision in storage. As a result, the activations and model weights are kept at their original precision while the inputs and outputs of matrix multiplication are dynamically masked to emulate reduced precision computation with a high-precision accumulator. All states are kept in their original precision, and all operations other than matrix multiplication are performed in their original precision. In Figure 4, we include a diagram indicating the precision of the states and computations in a Llama decoder block.
Figure 4: Diagram showing the precisions used in a Llama decoder block (best seen in color). The activations in the path of the residual connection are kept in FP32, as are the model weights and embeddings. The LayerNorm and RoPE (Su et al., 2024) layers use FP32 internally for their computations. The Flash Attention kernel uses BF16 with no reduction in precision. All other layers use reduced-precision matrix multiplication that emulates low-precision computation with a high-precision accumulator.
@@ 4.1 MS-AMP experiments
Figure 5: Comparison between MS-AMP FP8 (O1) and BF16 training on a subsample of the FineWeb Edu dataset using a Llama 120M model.
Figure 6: Training losses for GPT-2 training compared using exponential moving averages to better visualize the general trends. The blue curve indicates the training loss for the baseline BF16 training, while the red curve indicates MS-AMP level O1 training with the LM head excluded from FP8 quantization. The green curve shows the training loss for when the LM head included in the FP8 quantization.
We first analyze the effect of real-world FP8 training by applying MS-AMP (Peng et al., 2023) v0.4.0 level O1. All experiments are run on an H100 node to ensure hardware availability of FP8 with 8 GPUs. Two sets of experiments are performed. In the first, shown in Figure 5, we train a Llama 120M model on a subsample of the FineWeb Edu (Penedo et al., 2024) dataset. In the second, shown in Figure 6, we train a GPT-2 124M model from the nanoGPT (Karpathy, 2022) codebase using the OpenWebText (Gokaslan & Cohen, 2019) dataset.
From the two experiments, we come to the following conclusions. In Figure 5, where we train a Llama model on FineWeb Edu, a well-known “clean” dataset, we do not observe a difference between the convergence rates of MS-AMP O1 models and BF16 reference models. However, in Figure 6, we see a divergence between the MS-AMP O1 results and the BF16 baseline that does not close, even after applying six times the training iterations for the MS-AMP model at 120K training steps. The differences are much greater when the LM head is included in the quantization, indicating that parts of the model are highly sensitive to the quantization, requiring further investigation into model architecture choices.
These results indicate that the FP8 training scheme in MS-AMP may not converge to the same loss as BF16 training or requires more training steps, depending on factors such as data quality. This strengthens our case that FP8 training may introduce hidden instabilities that are not evident until stress tested against circumstances that were not considered in the original works proposing them.
LLM pretraining datasets in production environments, especially for customized use-cases, are usually much “dirtier” than those of popular open-source datasets. For example, extensive data filtering may not be an option for low-resource languages or for subfields such as medical or legal data. Also, for newer domains such as video or robotic motion, well-established metrics of data quality do not yet exist. Therefore, the finding that a method of FP8 training may work well only on “clean” data would be a consequential one for LLM training in production.
@@ 4.2 Bit reduction experiments
Figure 7: TinyLlama 120M models trained until loss divergence. E8M4 and E8M5 models are trained for 20K and 100K steps, respectively. The dotted black line in each figure indicates the loss landscape sharpness of the model.
We find that removing even a single exponent bit prevents training altogether, resulting in the model failing to progress with any learning using E7M7, confirming previous findings (Henry et al., 2019) that neural network training is more sensitive to exponent bits than mantissa bits. To analyze the cause, we conduct an ablation on the clamping mechanism by either removing only the inner or outer exponent range, as depicted in Figure 10 in the Appendix. We find that models with only their inner exponent ranges clamped train normally while models with only their outer exponent ranges clamped do not, indicating that the inability to represent large values is the cause of failure for E7M7. We therefore investigate the effects of removing mantissa bits for the remainder of our experiments.
4.3 Loss landscape sharpness
To further uncover the relationship between bit width and training robustness, we use Equation 2 to quantify the degree of training instability increase by measuring the loss landscape instability of Llama models. In Figure 7, we show Llama 120M models trained until their training losses diverge, as well as plotting the loss landscape sharpness values of the models in E8M4 and E8M5. Training required approximately 2 hours per 10,000 training steps on a node with 8 A100 GPUs.
Although the points of divergence are different for each model, we can see a general trend of increasing sharpness until the model diverges sharply, after which it cannot revert to its original training trajectory. This pattern is visible despite the large differences in training steps.
To verify that similar behavior is reproducible in larger models, we compare the training losses of Llama 7B models trained for 5,000 steps in Figure 8 and show the measured sharpness values for $\epsilon=5e\mathrm{-4}$ in Table 1. Results for other $\epsilon$ values are included in Table To FP8 and Back Again: Quantifying Reduced Precision Effects on LLM Training Stability of the Appendix and show a similar pattern. We apply early stopping at 5,000 training steps because training a Llama 7B model for 5K steps requires approximately one week on a single node with 8 A100 GPUs. These experiments show that loss divergence is visible in the E8M3 and E8M4 models, while it has yet to emerge in the E8M5 model. However, from Table 1, we can see that the loss landscape sharpness continues to increase for the E8M5 model, even though no signs of instability are yet visible in the training loss.
Figure 8: Llama 7B model training loss curves for different mantissa bits. The x-axis shows training steps, while the y-axis shows the training loss.
@@ 4.3 Loss landscape sharpness
The E8M3 and E8M4 models show much higher sharpness values, and both diverge early in training. In contrast, there is only a gradual increase in the loss-landscape sharpness for the E8M7 runs. Figures 7 and 8 show that models gradually increase in sharpness until a threshold level is reached. However, the exact threshold may differ depending on the configurations. These results suggest that models with fewer mantissa bits enter regions of ever greater instability during training, even when these instabilities are not visible in the loss curve. We believe that, in the future, such analysis of loss landscape sharpness can be used to identify when the model is at risk of training loss divergence.
To verify that our proposed loss landscape sharpness metric tracks the training instability when it decreases as well as when it increases, we show the loss landscape sharpness trend for the E8M3 model when training is reverted to BF16 in Table To FP8 and Back Again: Quantifying Reduced Precision Effects on LLM Training Stability of the Appendix. Supporting our claim that the loss landscape sharpness is an indicator of training instability, the sharpness value continues to decrease until it nears the level of BF16 training as shown in Table 1.
@@ 4.4 Robustness to learning rate changes
We further attempt to identify hidden instability in E8M5, which did not diverge during the initial training stages in Figure 7. Inspired by Wortsman et al. (2024), we analyze the robustness of Llama 120M models to changes in the learning rate by comparing training at BF16 with that for E8M5. As seen in Figure 9, the E8M5 training runs have more frequent loss spikes during training, especially when the learning rate is increased to $4e\mathrm{-3}$. Although no cases of loss divergence were found, we believe that the higher frequency of loss spikes indicates greater sharpness of the loss landscape, supporting our claim that training is more unstable for E8M5 even before loss divergence occurs.
@@ 5 Discussion
This work proposes quantitative evaluations and analyses of training instabilities when reducing floating-point precision. Our experiments have shown that MS-AMP, an existing open-source implementation of FP8 training, does not offer sufficient robustness to allow its cost-effective use. By emulating lower-precision mantissa bits in matrix multiplication, we further analyze the increasing instabilities arising from lower precision in LLM training.
Again, we clarify that we are not arguing that FP8 training is not viable. Indeed, we discuss the unique characteristics of DeepSeek V3 FP8 training in Section 2.1 and reproduce stable training using FP8 using the MS-AMP library in Section 4.1. The issue is that FP8 training causes a narrowing of the hyperparameter space where LLM training can occur stably and with equivalent performance as mixed-precision BF16 training. Because of this narrower hyperparameter space, more resources must be expended on identifying and honing techniques for preventing training collapse. Worse, there is simply no way of knowing if the FP8 training is performing competitively as BF16 without implementing a BF16 training run for comparison, which would negate the purpose of using FP8 for training in the first place. Even if FP8 training is practical for carefully selected hyperparameters under specific conditions, we assert that the costs of finding such conditions and the risks involved in training a less stable model outweigh the benefits of using FP8 for accelerated computation. Because of this risk, we propose methods to evaluate the stability and robustness of reduced-precision training, which is vital for FP8 or other reduced-precision training schemes to be viable for real-world LLM training.
From our experiments, several methods naturally suggest themselves as possible stabilization techniques. First, the initial stages of training could be conducted in higher precision, similar to how smaller batch sizes may be used during the initial stages of training as in Keskar et al. (2017). Increasing the precision when the loss landscape becomes too sharp may also provide a tradeoff between training speed and stability. Second, the more sensitive layers may be kept at high precision, while only the less sensitive layers are computed with reduced precision. For example, in Figure 11, we found that removing masking from the LM head of a Llama model was sufficient to enable E7M7 training to progress, although the resulting model was unstable. For the instability in GPT models shown in Figure 1, we found that increasing the precision of the first two decoder blocks to TF32 was sufficient to prevent loss divergence. However, as such compensatory techniques depend on the model architecture, training data, and other aspects of the training environment, we leave their investigation to future work.
@@ 6 Limitations
A limitation of this work is that it focuses on the initial stages of pre-training when many instabilities are known to arise only later in training (Fishman et al., 2024; Bekman, 2023). For example, Wortsman et al. (2024) show that the logits of the outputs diverge from zero only at the later stages of training, and Fishman et al. (2024) show that correlations between the SwiGLU layers only begin to increase after 200B tokens. To this, we argue that our studies likely underestimate the instabilities that FP8 or other reduced precision training schemes will face, further strengthening our argument that currently available FP8 training methods are too unstable to be profitably utilized for LLM training in their current form and that much more extensive evaluations are required to stress-test them.
Second, despite finding that the exponent bits are of greater importance to LLM training than mantissa bits, we were unable to experiment by increasing the number of exponent bits. This was because matrix multiplication in FP64 is over an order of magnitude slower than BF16 on A100 and H100 GPUs when using tensor cores. Experiments using representations such as E11M4, created by removing 48 mantissa bits from FP64, may be illuminating, but we found it impractical to train models with a greater number of exponent bits due to hardware limitations.
Finally, our experiments are limited in that they only the training loss is used as an evaluation metric instead of real-world natural language tasks such as MMLU (Hendrycks et al., 2021) scores. However, while lower perplexity does not guarantee superior performance on downstream tasks, we believe that training loss divergence is sufficient to indicate training failure.
@@ 7 Conclusion
We demonstrate that the training stability of LLMs decreases incrementally with the reduction of floating-point bit widths used for training models of up to 7B parameters. Using our proposed loss landscape sharpness metric, we measure the gradual narrowing of the hyperparameter space where stable training is possible.
In future works proposing reduced-precision training, we hope that our work inspires analyses of the robustness of newly proposed training methods not only on the well trodden paths with well-established guidelines, but also on less explored conditions where many of the initial assumptions may not hold. The massive power and capital consumption of LLM training means that much is at stake on this issue.